# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

### v1

@@ Abstract.
Rényi divergences play a pivotal role in information theory, statistics, and machine learning. While several estimators of these divergences have been proposed in the literature with their consistency properties established and minimax convergence rates quantified, existing accounts of probabilistic bounds governing the estimation error are premature. Here, we make progress in this regard by establishing exponential deviation inequalities for smoothed plug-in estimators and neural estimators by relating the error to an appropriate empirical process and leveraging tools from empirical process theory. In particular, our approach does not require the underlying distributions to be compactly supported or have densities bounded away from zero, an assumption prevalent in existing results. The deviation inequality also leads to a one-sided concentration bound from the expectation, which is useful in random-coding arguments over continuous alphabets in information theory with potential applications to physical-layer security. As another concrete application, we consider a hypothesis testing framework for auditing Rényi differential privacy using the neural estimator as a test statistic and obtain non-asymptotic performance guarantees for such a test.
@@ 1. Introduction
Rényi divergences play a fundamental role in information theory, statistics, machine learning, and related fields. These divergences as well as information measures derived from them, such as mutual information and entropy[54], characterize fundamental limits in information-theoretic coding problems. For instance, they characterize the error-exponents in hypothesis testing, the capacity of noisy channels, ultimate limits in data compression and rate-distortion theory, to name a few (see e.g., [11, 16]). Further, they appear as discrepancy measures or as optimization metrics in applications. Rényi divergence is a functional of the underlying probability distributions, which are often unknown in practice. In such scenarios, the problem of estimating Rényi divergences from samples arises, for which various estimators have been proposed in the literature, such as those based on plug-in, kernel, and variational methods (see Section 1.3 below). While consistency and minimax error rates for several of these estimators have been established, an understanding of the probability law governing the estimation error, which is pivotal for a complete statistical account of performance, is limited.
In this work, we make progress in this regard by establishing exponential deviation inequalities for two classes of Rényi divergence estimators, namely Gaussian-smoothed plug-in estimator and neural estimator. To this end, we utilize known variational expressions for these divergences (see (1) and (2) below)[19, 50, 10]. Our approach relies on relating the deviation to the suprema of empirical processes indexed by suitable function classes, and using concentration inequalities for the suprema to arrive at the desired conclusion. Our results go beyond the framework of discrete alphabets and compactly supported distributions with densities bounded away from zero, as required in existing works (see e.g., [58, 59, 65]). Specifically, we obtain deviation inequalities for Gaussian-smoothed empirical measures (based on samples) which satisfy neither of these criteria, under the assumption that the unsmoothed distribution is compactly supported or sub-Gaussian. Further, under an additional assumption that the densities have bounded ratios and belong to an appropriate Lipschitz class, we are also able to treat the deviation of the estimate from the unsmoothed value. We also obtain exponential inequalities for neural estimators of Rényi divergences and discuss its application to obtaining the first complete non-asymptotic performance guarantee for auditing Rényi differential privacy (DP) [33, 18] using the neural estimator as a test statistic.
@@ 1.1. Estimation of Rényi divergences
Our approach to deriving the deviation inequality for Gaussian smoothed plug-in estimator relies on identifying primitive conditions on the distributions which allow the estimation error to be controlled by the suprema of empirical processes with good concentration properties. In this regard, we show that when $\mu$ and $\nu$ are compactly supported or sub-Gaussian, it is possible to relate the desired deviation to suprema of empirical processes indexed by a smooth function class with an integrable envelope [^3] An envelope of a class $\mathcal{F}$ of functions on $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a function $F:\mathcal{X}\rightarrow[0,\infty]$ such that $\sup_{f\in\mathcal{F}}\left|f(x)\right|\leq F(x)$ for all $x\in\mathcal{X}$.(in a suitable Orlicz sense). The integrability of the envelope then translates to a concentration property for the suprema as quantified by a Fuk-Nagaev type inequality[24, 2]. In order to apply the latter, we require an upper bound on the covering entropy (see Definition 3) of a Gaussian-smoothed function class with an envelope $F(x)$ that scales at most as some power of $\left\|x\right\|$. As for neural estimators, we leverage a version of Talagrand’s concentration inequality given in [12] to obtain the desired exponential inequality.
@@ 1.2. Contributions
    $\mathbb{P}\left(\left|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\geq\Omega_{\alpha,r,d,\sigma}\big{(}n^{-\frac{1}{2}}+z\big{)}\right)\leq 2e^{-\left(nz^{2}\wedge nz\right)}.$ (4)
@@
Exponential concentration inequalities combined with random coding- or expurgation-based arguments are useful in information-theoretic security to show existence of coding schemes satisfying secrecy criteria quantified in terms of a divergence measure (see e.g., [44, 25, 63, 64]). The deviation inequalities (3)-(4) are continuous alphabet analogues of such inequalities for Rényi divergences, which, to the best of our knowledge, are known currently only for finite alphabets. Hence, we expect that these inequalities will have potential applications to secure and covert communications over a Gaussian wiretap channel[43]. As a concrete application of (5), we consider a hypothesis testing based framework for auditing Rényi DP and obtain non-asymptotic performance guarantees for this test. While this application has been considered before in [18] and [67], the auditing guarantees were either partial or only valid asymptotically. Here, we address these shortcomings thus making the auditing framework more appealing in practice.
@@ 1.3. Related work
The literature on estimation of information and divergence measures is extensive (see e.g., [77] for a survey). Several estimators of such measures have been proposed and their statistical properties analyzed. In this regard, consistency, convergence rates, and minimax rates of various estimators based on plug-in, kernel, k-nearest neighbour, and transform methods have been studied in [39, 72, 4, 52, 40, 78, 15, 42, 53, 50, 73, 69, 36, 34, 60, 80, 74, 51, 49, 13, 14, 9, 8, 28, 1, 26, 31, 46, 30, 82, 56, 70, 71, 32] (see also references therein). Further statistical properties such as the asymptotic distribution of the estimation error in the flavor of the classical central limit theorem have also been characterized, see e.g. [57, 36, 48, 38, 27, 66, 61, 67, 62]. Consistency of neural estimators of mutual information and Rényi divergences have been explored in [7] and [10], respectively, while minimax convergence rate for neural estimation of $f$-divergences over a sufficiently regular class of distributions has been established in [68, 65].
Probabilistic tail inequalities for kernel-based estimators of Rényi divergences and neural estimators of KL divergence for compactly supported distributions have been derived in [58, 59, 65]. For plug-in estimators of KL divergence, Chernoff-type concentration bounds have been studied in [29] for the finite alphabet setting, improving upon earlier works [3, 45]. However, all these results are limited to distributions with compact support and smooth densities bounded away from zero, which we aim to generalize upon.
@@ 2.1. Notation
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a sufficiently rich probability space on which all random variables are defined. Let $\rho$ be a sigma-finite positive measure on the Borel sigma algebra $\mathcal{B}(\mathbb{R}^{d})$ and $\mathcal{P}(\mathbb{R}^{d})$ be the set of Borel probability measures. For $\mu\in\mathcal{P}(\mathbb{R}^{d})$ such that $\mu\ll\rho$, we use $p_{\mu}=\frac{d\mu}{d\rho}$ to denote its density (Radon-Nikodym derivative), w.r.t. $\rho$. $\mu^{\otimes n}$ stands for the $n$-fold product measure, and $\delta_{x}$ represents the Dirac measure at $x$. $\mathbbm{1}_{\mathcal{E}}$ denotes the indicator of set $\mathcal{E}$. $\mu*\nu$ denotes the convolution of $\mu$ and $\nu$; likewise, fgf*g represents convolution of two measurable functions $f,g$. $X\sim\mu$ stands for random variable $X$ distributed according to $\mu$. The expectation of $X\sim\mu$, whenever it exists, will be denoted by $\mathbb{E}_{\mu}[X]$. We write $\gamma_{\sigma}=N(0,\sigma^{2}I_{d})$ for the centered Gaussian distribution on $\mathbb{R}^{d}$ with covariance matrix $\sigma^{2}I_{d}$, and use $\varphi_{\sigma}(x)=(2\pi\sigma^{2})^{-d/2}e^{-\left\|x\right\|^{2}/(2\sigma^{2})}$ for the corresponding density. Inequalities between random variables are interpreted to hold almost surely or surely, which will be evident from the context.
@@ 2.2. Definitions
Examples of Orlicz functions include $\hat{\psi}_{q}(z)=z^{q}$ and $\psi_{q}(z)=e^{z^{q}}-1$, $z\in\mathbb{R}$, for $q\geq 1$; in particular, $\psi_{2}$ corresponds to the sub-Gaussian class defined next.
@@
We also require the following technical assumption on certain function classes of interest.
@@ 3.1. Deviation Inequalities for Smoothed Plug-in Estimators
    where $\hat{\xi}_{d,L,\sigma}$, $\check{\xi}_{d,L,\sigma}$ and $\hat{\zeta}_{n,d,p,\tau}(z)$ are given in (42b), (42c) and (43), respectively.
The proof of Theorem 1 is given in Section 4.1 and relies on a more general result (Lemma 2) stated below. At a high level, the proof relies on relating the desired deviation to suprema of empirical processes indexed by appropriate function classes and leveraging powerful concentration inequalities from empirical process theory. More specifically, (8) utilizes Talagrand’s concentration inequality for bounded function classes given in [12, Theorem 2.3] (see Theorem 6 below), while (9) relies on a Fuk-Nagaev type inequality given in [2, Theorem 4] (see Theorem 7 below) that applies to unbounded function classes. To apply these results, we also obtain an upper bound on the covering entropy of Gaussian-smoothed bounded function class (Lemma 3) by upper bounding the Hölder norm of such functions and utilizing a refinement of the covering entropy estimate for the Hölder class, which specifies the dependence on dimension and smoothness parameter in the prefactor.
@@
by convexity of KL divergence (see e.g., [76, Theorem 11]), Lemma 2 also yields a deviation inequality for the upper tail, i.e., (8) holds with $\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|$ replaced by $\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$. Similar claim also holds for (9). Further, an inspection of the proof shows that (8) also holds when $\hat{\mu}_{n}*\gamma_{\sigma}$ or $\hat{\nu}_{n}*\gamma_{\sigma}$ (or both) are set to $\mu*\gamma_{\sigma}$ and $\nu*\gamma_{\sigma}$, respectively.
@@
$\mathcal{F}_{\sigma}(b)$ $:=\{f*\varphi_{\sigma}:\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|),~\forall~x\in\mathbb{R}^{d}\},$
$\mathcal{E}_{\sigma}(b)$ $:=\{e^{f}*\varphi_{\sigma}:\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|),~\forall~x\in\mathbb{R}^{d}\}.$
@@
The proof of Theorem 2 is given in Section 4.2 and follows by replacing Talagrand’s inequality in the proof of Lemma 2 by a statistical version involving symmetrized Rademacher expectations given in [6, Theorem 2.1]. In fact, a more general claim can be proven as stated below, which could be of independent interest.
@@
[…]
$Z_{n}$ $:=\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|,$ (11a)
$\mbox{and }Z^{(\varepsilon)}_{n}$ $:=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right].$ (11b)
@@
[…]
Lemma 1 can be considered as a generalization of [6, Theorem 2.1] to an unbounded function class whose envelope is integrable in a suitable Orlicz sense. The proof is given in Section 5.3. The essential idea is based on truncation of $\mathcal{F}_{n}$ to a bounded class of functions $\mathcal{F}_{n}(T_{n})=\left\{f\mathbbm{1}_{F_{n}\leq T_{n}}:f\in\mathcal{F}_{n}\right\}$, applying [6, Theorem 2.1] to the truncated class, invoking the Hoffmann-Jørgensen inequality (see e.g. [41, Proposition 6.8]) to control the remainder by selecting an appropriate $T_{n}$, and showing that $\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}(T_{n})}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right]\leq Z^{(\varepsilon)}_{n}$.
@@
This follows by applying Markov’s inequality instead of (91) in the proof of Lemma 1, which yields
$\mathbb{P}\left(Z_{n}^{\prime\prime}\geq\frac{z}{4}\right)\leq 4^{p}\frac{\mathbb{E}\big{[}Z_{n}^{\prime\prime p}\big{]}}{z^{p}}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{2(16)^{p}\mathbb{E}\left[M_{n}^{p}\right]}{z^{p}},$
where $(a)$ is via an application of Hoffmann-Jørgensen inequality given in [41, Proposition 6.8]:
$\mathbb{E}\big{[}Z_{n}^{\prime\prime p}\big{]}\leq 2(4^{p})\mathbb{E}\left[\max_{1\leq i\leq n}F_{n}^{p}\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right]\leq 2(4^{p})\mathbb{E}\left[M_{n}^{p}\right].$
@@
[…]

### v3

@@ Abstract.
Rényi divergences play a pivotal role in information theory, statistics, and machine learning. While several estimators of these divergences have been proposed in the literature with their consistency properties established and minimax convergence rates quantified, existing accounts of probabilistic bounds governing the estimation error are relatively underdeveloped. Here, we make progress in this regard by establishing exponential deviation inequalities for smoothed plug-in estimators and neural estimators by relating the error to an appropriate empirical process and leveraging tools from empirical process theory. In particular, our approach does not require the underlying distributions to be compactly supported or have densities bounded away from zero, an assumption prevalent in existing results. The deviation inequality also leads to a one-sided concentration bound from the expectation, which is useful in random-coding arguments over continuous alphabets in information theory with potential applications to physical-layer security. As another concrete application, we consider a hypothesis testing framework for auditing Rényi differential privacy using the neural estimator as a test statistic and obtain non-asymptotic performance guarantees for such a test.
@@
S. Sreekumar is supported in parts by the Centre National de la Recherche Scientifique (CNRS) funding and the CentraleSupélec-L2S funding. K. Kato is partially supported by NSF grant DMS-2413405.
@@ 1. Introduction
Rényi divergences [64] play a fundamental role in information theory, statistics, machine learning, and related fields. These divergences as well as information measures derived from them, such as mutual information and entropy, characterize fundamental limits in information-theoretic coding problems. For instance, they characterize the error-exponents in hypothesis testing, the capacity of noisy channels, ultimate limits in data compression and rate-distortion theory, to name a few (see e.g., [19]). Further, they appear as discrepancy measures or as optimization metrics in applications. Rényi divergence is a functional of the underlying probability distributions, which are often unknown in practice. In such scenarios, the problem of estimating Rényi divergences from samples arises, for which various estimators have been proposed in the literature, such as those based on plug-in, kernel, and variational methods (see Section 1.3 below). While consistency and minimax error rates for several of these estimators have been established, an understanding of the probability law governing the estimation error, which is pivotal for a complete statistical account of performance, is limited.
In this work, we make progress in this regard by establishing exponential deviation inequalities for two classes of Rényi divergence estimators, namely Gaussian-smoothed plug-in estimator and neural estimator. To this end, we utilize known variational expressions for these divergences (see (1) and (2) below) [24, 58, 13]. Our approach relies on relating the deviation to the suprema of empirical processes indexed by suitable function classes, and using concentration inequalities for the suprema to arrive at the desired conclusion. Our results go beyond the framework of discrete alphabets and compactly supported distributions with densities bounded away from zero, as required in existing works (see e.g., [69, 70, 75]). Specifically, we obtain deviation inequalities for Gaussian-smoothed empirical measures (based on samples) which satisfy neither of these criteria, under the assumption that the unsmoothed distribution is compactly supported or sub-Gaussian. Further, under an additional assumption that the densities have bounded ratios and belong to an appropriate Lipschitz class, we are also able to treat the deviation of the estimate from the unsmoothed value. We also obtain exponential inequalities for neural estimators of Rényi divergences and discuss their applications to obtaining the first complete non-asymptotic performance guarantee for auditing Rényi differential privacy (DP) [39, 23] using the neural estimator as a test statistic.
@@ 1.1. Estimation of Rényi divergences
Our approach to deriving the deviation inequality for Gaussian-smoothed plug-in estimator relies on identifying primitive conditions on the distributions which allow the estimation error to be controlled by the suprema of empirical processes with good concentration properties. In this regard, we show that when $\mu$ and $\nu$ satisfy certain regularity conditions, which are in particular satisfied in the compactly supported or sub-Gaussian case, it is possible to relate the desired deviation to suprema of empirical processes indexed by a smooth function class with an integrable envelope [^3] An envelope of a class $\mathcal{F}$ of functions on $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a function $F:\mathcal{X}\rightarrow[0,\infty]$ such that $\sup_{f\in\mathcal{F}}\left|f(x)\right|\leq F(x)$ for all $x\in\mathcal{X}$.. The integrability of the envelope then translates to a concentration property for the suprema as quantified by a version of Talagrand’s concentration inequality given in [15] (or a Fuk-Nagaev type inequality[29, 2, 3] in the general setting). In order to apply these inequalities, we require an upper bound on the covering entropy (see Definition 3) of a Gaussian-smoothed function class with an envelope $F(x)$ that scales at most as some power of $\left\|x\right\|$. As for neural estimators, we again leverage Talagrand’s concentration inequality to obtain the desired exponential inequality.
@@ 1.2. Contributions
    $\mathbb{P}\left(\left|\mathsf{D}_{\alpha}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|\geq\Omega_{r,d,\sigma,\alpha}\big{(}n^{-\frac{1}{2}}+z\big{)}\right)\leq 2e^{-\left(nz^{2}\wedge nz\right)}.$ (4)
@@
As our key technical contributions, we derive a more general lemma (Lemma 2) that provides deviation inequalities for smoothed plug-in estimator of KL divergence under certain regularity conditions involving the smoothing kernel and underlying probability measures. Analogous lemma can also be stated for Réyi divergences, however, we omit this to avoid repetition. A key tool required to apply Lemma 2 to prove Theorem 1 and 3 is a bound on the covering entropy of a Gaussian-smoothed function class whose envelope scales as an affine or quadratic function of $\left\|x\right\|_{2}$. To this end, we obtain a uniform upper bound (Lemma 3) on the Hölder norm of functions in this class, and use the covering entropy estimate for Hölder class given in [83, Theorem 2.7.1]. In fact, we apply a slightly refined version of this result post quantifying the dependence of dimension and support (see Theorem 8). Leveraging [7, Theorem 2.1], we also provide a statistical version (Theorem 2) of (3) in terms of Rademacher complexity of the Gaussian-smoothed function class, which is in general tighter than (3). We also prove a generalization of [7, Theorem 2.1] to function classes with unbounded envelope (Lemma 1), which is a statistical version of Fuk-Nagaev type inequality given in [2, Theorem 4] and could be of independent interest.
Exponential concentration inequalities combined with random coding- or expurgation-based arguments are useful in information-theoretic security to show existence of coding schemes satisfying secrecy criteria quantified in terms of a divergence measure (see e.g., [54, 31, 73, 74]). The deviation inequalities (3)-(4) are continuous alphabet analogues of such inequalities for Rényi divergences, which, to the best of our knowledge, are known currently only for finite alphabets or compactly supported distributions. Hence, we expect that these inequalities will have potential applications to secure and covert communications over a Gaussian wiretap channel[52]. As a concrete application of (5), we consider a hypothesis testing based framework for auditing Rényi DP and obtain non-asymptotic performance guarantees for this test. While this application has been considered before in [23] and [76], the auditing guarantees were either partial or only valid asymptotically. Here, we address these shortcomings thus making the auditing framework more appealing in practice.
@@ 1.3. Related Work
The literature on estimation of information and divergence measures is extensive, hence we will only focus on the literature relevant to Rényi divergence estimators, which includes KL divergence, mutual information and entropy estimators as a special case. Several estimators of these measures have been proposed and their statistical properties analyzed. In this regard, consistency, convergence rates, and minimax rates of various estimators based on plug-in, kernel and k-nearest neighbour methods have been studied in [80, 20, 5, 61, 68, 47, 33, 85, 18, 51, 62, 35, 58, 60, 81, 63, 77, 48, 42, 40, 71, 87, 45, 59, 21, 82, 30, 57, 16, 86, 10, 1, 32, 36, 37, 89, 17, 14, 66, 9, 78, 79, 38] (see also references therein). Further statistical properties such as the asymptotic distribution of the estimation error in the flavor of the classical central limit theorem have also been characterized, see e.g. [42, 57, 10, 9, 76, 72]. While many of the above references assume compactly supported distributions with smooth densities bounded away from zero, some of the results hold more generally. Notably, [10] and [9] show that weighted average of Kozachenko–Leonenko estimators [46] are minimax-rate optimal for entropy and divergence estimation, respectively, over a general class of locally smooth distributions (which includes multi-variate Gaussians) under the squared error criterion. For the same error criterion, [37] establish minimax rate optimality of a kernel-based entropy estimator over a Lipschitz class of densities supported on $0,1]^{d}$. Consistency of a neural estimator of Rényi divergence has been explored in [13], while minimax convergence rate for neural estimation of KL divergence over a sufficiently regular class of distributions has been established in [75].
Probabilistic tail inequalities for kernel-based estimators of Rényi divergences and mutual information as well as neural estimators of KL divergence for compactly supported distributions have been derived in [69, 70, 53, 75]. For plug-in estimators of KL divergence, Chernoff-type concentration bounds have been studied in [34] for the finite alphabet setting, building upon earlier works [4, 55]. However, all these results are limited to distributions with compact support and smooth densities bounded away from zero, which we aim to generalize upon.
@@ 2.1. Notation
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a sufficiently rich probability space on which all random variables are defined. Let $\rho$ be a sigma-finite positive measure on the Borel sigma algebra $\mathcal{B}(\mathbb{R}^{d})$ and $\mathcal{P}(\mathbb{R}^{d})$ be the set of Borel probability measures. For $\mu\in\mathcal{P}(\mathbb{R}^{d})$ such that $\mu\ll\rho$, we use $p_{\mu}=\frac{d\mu}{d\rho}$ to denote its density (Radon-Nikodym derivative), w.r.t. $\rho$. $\mu^{\otimes n}$ stands for the $n$-fold product measure, and $\delta_{x}$ represents the Dirac measure at $x$. $\mathbbm{1}_{\mathcal{E}}$ denotes the indicator of set $\mathcal{E}$. $\mu*\nu$ denotes the convolution of $\mu$ and $\nu$; likewise, fgf*g represents convolution of two measurable functions $f,g$. $X\sim\mu$ stands for random variable $X$ distributed according to $\mu$. The expectation of $X\sim\mu$, whenever it exists, will be denoted by $\mathbb{E}_{\mu}[X]$. We write $\gamma_{\sigma}=N(0,\sigma^{2}I_{d})$ for the centered Gaussian distribution on $\mathbb{R}^{d}$ with covariance matrix $\sigma^{2}I_{d}$, and use $\varphi_{\sigma}(x)=(2\pi\sigma^{2})^{-d/2}e^{-\left\|x\right\|^{2}/(2\sigma^{2})}$ for the corresponding density. Inequalities between random variables are interpreted to hold almost surely (a.s.) or surely, which will be evident from the context.
@@ 2.2. Definitions
Examples of Orlicz functions include $\hat{\psi}_{p}(z)=z^{p}$ and $\psi_{p}(z)=e^{z^{p}}-1$, $z\in\mathbb{R}$, for $p\geq 1$; in particular, $\psi_{2}$ corresponds to the sub-Gaussian class defined next.
@@
We further require the following technical assumption on certain function classes of interest.
@@ 3.1. Deviation Inequalities for Smoothed Plug-in Estimators
    where
    $\hat{\zeta}_{n,d,p,\tau}(z):=n^{\frac{d+12-8\tau}{d+12}}z^{2}\wedge n^{\frac{d+12-4\tau}{d+12}}z\wedge\frac{n^{\frac{p\tau}{d+12}}}{\log(n+1)},$ (10)
    and $\hat{\xi}_{d,L,\sigma}$, $\check{\xi}_{d,L,\sigma}$ are given in (41).
The proof of Theorem 1 is given in Section 4.1 and relies on a more general result (Lemma 2) stated below. At a high level, the proof relies on relating the desired deviation to suprema of empirical processes indexed by appropriate function classes and leveraging concentration inequalities from empirical process theory. More specifically, Lemma 2 relies on a Fuk-Nagaev type inequality given in [3, Theorem 2] (see Theorem 7 below) that applies to an unbounded function class with an integrable envelope. We also obtain an upper bound on the covering entropy of Gaussian-smoothed bounded function class (Lemma 3) by upper bounding the Hölder norm of such functions and utilizing a refinement of the covering entropy estimate for the Hölder class, which specifies the dependence on dimension and smoothness parameter in the prefactor.
@@
Observe that (8) holds for compactly supported $\mu,\nu$ in the two-sample setting (where both $\mu$ and $\nu$ are approximated), while (9) holds more generally for sub-Gaussian $\mu,\nu$, although limited to the one-sample setting (where only $\mu$ is approximated). The latter restriction arises from the limitation that in order to upper bound the desired deviation in terms of empirical process suprema and invoke Talagrand’s inequality or Fuk-Nagaev type inequality, we require $f_{\hat{\mu}_{n}*\varphi_{\sigma},\hat{\nu}_{n}*\varphi_{\sigma}}^{\star}$ to belong a.s. to a function class with an (Orlicz) integrable envelope. While this is always possible when $\mu$ and $\nu$ are compactly supported, this is no longer true when $\mu,\nu$ are sub-Gaussian in general. In the one-sample scenario, this limitation can be overcome by considering the event $E_{n}(t_{n}):=\left\{\max_{1\leq i\leq n}\left\|X_{i}\right\|\leq t_{n}\right\}$ to reduce to the compactly supported case, and bounding the probability of the complement event using the sub-Gaussianity of $\mu$. Then, optimizing the truncation level $t_{n}$ results in (9), with the third factor within the minimum in the RHS of (10) corresponding to the probability of $E_{n}(t_{n})$ complement. Although this approach is also possible in principle for the two-sample case, it does not lead to exponential inequalities since the relevant norm of the envelope scales exponentially in the truncation level.
@@
by convexity of KL divergence (see e.g., [84, Theorem 11]), Lemma 2 also yields a deviation inequality for the upper tail, i.e., (8) holds with $\left|\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\right|$ replaced by $\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)-\mathbb{E}\left[\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)\right]$. Similar claim also holds for (9). Further, an inspection of the proof shows that (8) also holds when $\hat{\mu}_{n}*\gamma_{\sigma}$ or $\hat{\nu}_{n}*\gamma_{\sigma}$ (or both) are set to $\mu*\gamma_{\sigma}$ and $\nu*\gamma_{\sigma}$, respectively.
@@
$\mathcal{F}_{\varphi_{\sigma}}(b)$ $:=\{f*\varphi_{\sigma}:\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|),~\forall~x\in\mathbb{R}^{d}\},$
$\mathcal{E}_{\varphi_{\sigma}}(b)$ $:=\{e^{f}*\varphi_{\sigma}:\left|f(x)\right|\leq 0.5b(1+\left\|x\right\|),~\forall~x\in\mathbb{R}^{d}\}.$
@@
The proof of Theorem 2 is given in Section 4.2 and follows by replacing Talagrand’s inequality in the proof of Lemma 2 by a statistical version involving symmetrized Rademacher expectations given in [7, Theorem 2.1]. In fact, a more general claim can be proven as stated below, which could be of independent interest.
@@
Let $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ be such that $\mathcal{F}_{n}$ satisfies Assumption 1 with envelope $F_{n}$, and suppose $\left\|F_{n}\right\|_{\psi_{q}}<\infty$ for some $q\in(0,1]$ and all $n\in\mathbb{N}$. For $X^{n}\sim\mu^{\otimes n}$ and i.i.d. Rademacher variables $\{\varepsilon_{i}\}_{i=1}^{n}$ independent of $X^{n}$, set $M_{n}(X^{n}):=\max_{1\leq i\leq n}F_{n}(X_{i})$, $\overline{M_{n}}:=\mathbb{E}\big{[}\max_{1\leq i\leq n}F_{n}(X_{i})\big{]}$,
$Z_{n}$ $:=\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|\quad\mbox{and}\quad Z^{(\varepsilon)}_{n}:=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right].$ (12)
@@
$\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}(8\overline{M_{n}})}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right].$
Lemma 1 can be considered as a generalization of [7, Theorem 2.1] to an unbounded function class whose envelope is integrable in a suitable Orlicz sense. The proof is given in Section 5.3. The essential idea is based on truncation of $\mathcal{F}_{n}$ to a bounded class of functions $\mathcal{F}_{n}(T_{n})=\left\{f\mathbbm{1}_{F_{n}\leq T_{n}}:f\in\mathcal{F}_{n}\right\}$, applying [7, Theorem 2.1] to the truncated class, invoking the Hoffmann-Jørgensen inequality (see e.g. [50, Proposition 6.8]) to control the remainder by selecting an appropriate $T_{n}$, and showing that $\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}_{n}(T_{n})}\left|\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|\right]\leq Z^{(\varepsilon)}_{n}$.
@@
This follows by applying Markov’s inequality (in place of (85)) in the proof of Lemma 1, which yields
$\mathbb{P}\left(Z_{n}^{\prime\prime}\geq\frac{z}{4}\right)\leq 4^{p}\frac{\mathbb{E}\big{[}Z_{n}^{\prime\prime p}\big{]}}{z^{p}}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{2(16)^{p}\mathbb{E}\left[\big{(}M_{n}(X^{n})\big{)}^{p}\right]}{z^{p}},$
where $(a)$ is via an application of Hoffmann-Jørgensen inequality given in [50, Proposition 6.8]:
$\mathbb{E}\big{[}Z_{n}^{\prime\prime p}\big{]}\leq 2(4^{p})\mathbb{E}\left[\max_{1\leq i\leq n}F_{n}^{p}\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right]\leq 2(4^{p})\mathbb{E}\left[\big{(}M_{n}(X^{n})\big{)}^{p}\right].$
@@
Lemma 1 is useful for constructing confidence intervals for $Z_{n}$ when $\mu$ is unknown, provided $Z^{(\varepsilon)}_{n}$ can be computed and upper bounds on $\left\|F_{n}(X)\right\|_{\psi_{q}}$ are known (since $\left\|F_{n}\right\|_{2,\mu}$, $\overline{M_{n}}$ and $\left\|M_{n}(X^{n})\right\|_{\psi_{q}}$ can be upper bounded in terms of this).